### v2

@@ 2.5 Axioms
While $\text{C}^{2}$ captures numerous graph properties, it falls short in expressing certain important characteristics, such as trees, forests, and strongly connected orientations. In fact, these structures cannot be finitely axiomatized by first-order logic.
@@
An axiom is a special constraint on the interpretation of a binary predicate $R$ that $G(R)$ should be a certain combinatorial structure. We often write an axiom as $axiom(R)$ where $axiom$ is an identifier of the axiom. For example, in what follows, we will discuss the tree axiom $tree(R)$, which requires $G(R)$ to be a tree, and the forest axiom $forest(R)$, which requires $G(R)$ to be a forest. For better illustration, we treat axioms as atomic formulas in a sentence, and use $\Psi\land axiom(R)$ to denote the sentence $\Psi$ with the axiom $axiom(R)$ conjoined. For instance, $tree(E)\land\forall x\left(Leaf(x)\leftrightarrow\exists_{=1}y\ E(x,y)\right)$ is a sentence that expresses the concept of trees with a leaf predicate $Leaf/1$.
@@ 2.6 Graph Polynomial
Tutte Polynomial. The Tutte polynomial [Tut54] is one of the most well-known graph polynomials for undirected graphs. Consider an undirected graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges. The Tutte polynomial for the graph $G$ is a polynomial in two variables $x$ and $y$ defined as
@@
The Tutte polynomial holds significant combinatorial meaning. Evaluating the polynomial at specific points reveals various combinatorial properties. For example, $T_{G}(1,1)$ counts the number of spanning forests of $G$ (specifically, the number of spanning trees if $G$ is connected), and $T_{G}(0,2)$ counts the strongly connected orientations of $G$ [Ver80]. Substituting a concrete value for only one variable yields another graph polynomial of $G$. For example, $T_{G}(x,0)$ corresponds to the chromatic polynomial [Rea68] of $G$; $T_{G}(0,y)$ corresponds to the flow polynomial [DSS08] of $G$.
The computation of the Tutte polynomial for an arbitrary graph is known to be computationally challenging. Even the evaluation of a specific point of the polynomial is known to be #$\mathsf{\#P}$-hard, with only a limited number of points being computationally tractable [JVW90].
Directed Chromatic Polynomial. The idea of the chromatic polynomial of a directed graph appeared in [Sta70], although it was originally stated in terms of ordered sets. Given a directed graph $D=(V,E)$ where $n=|V|$, let $\chi_{D}(x)$ be the number of ways of coloring each vertex in $D$ by one of the colors $\{1,2,\cdots,x\}$ such that if there is an edge from $u$ to $v$, the color of $u$ is smaller than the color of $v$. We have the following lemma to state that $\chi_{D}(x)$ is a polynomial of $x$.
Lemma 2.13 ([Sta70]).
@@
Lemma 2.16 ([Sta70, AB20]).
@@
$\chi_{D}(x)$ =$=\begin{cases}(-1)^{n}\bar{\chi}_{D}(-x),&D\text{ is acyclic,}\\ 0,&\text{otherwise,}\end{cases}$
@@ 3.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w,\overline{w}$ be two weighting functions, and $R$ be a distinguished binary relation. The $n$-th WCP $f_{n}(u)$ of $\Psi$ for the relation $R$ is the unique univariate polynomial which satisfies:
@@
For the $n$-th WCP of the sentence $\Psi$, the weighting functions $w$ and $\overline{w}$ and the binary relation $R$, it holds that
@@
The $n$-th WCP $f_{n}(u;\Psi,w,\overline{w},R)$ is a univariate polynomial of degree at most $n$.
@@
Since $cc(\mu_{R})\leq n$, it follows immediately from Proposition 3.2. ∎
@@ 3.2 Fast Calculation
Note that by definition, evaluating the polynomial at a specific positive integer is a WFOMC query. However, such WFOMC is not readily solvable by existing algorithms (e.g., [BdBGS15, vBK21]) even for the $\text{FO}^{2}$ fragment. In general, the runtime of these algorithms is exponential in the length of the sentence, while the length of the sentence $\Psi_{R,u}$ in the WFOMC query can be as large as the domain size $n$, hence these algorithms would not lead to polynomial runtime in the domain size. In particular, the algorithm described in Section 2.4 is not sufficient in this case either, as there would be $O(2^{n})$ 1-types and, consequently, an exponential time data complexity. To this end, we adapt this algorithm into a dynamic programming (DP) style algorithm that exploits the structure of $\Psi_{R,u}$ to achieve an efficient calculation of the polynomials for $\text{FO}^{2}$ sentences, which then applies to $\text{C}^{2}$ sentences with cardinality constraints by techniques mentioned in Section 2.4.
Let $\Psi$ and $R$ be the input $\text{FO}^{2}$ sentence and the binary relation that WCP is defined on. Consider the point evaluation of the WCP at $u$. Let $C_{1}(x),C_{2}(x),\cdots,\linebreak C_{L}(x)$ be the valid 1-types  [^7] We call a 1-type valid if it is satisfiable in some model of the sentence. of $\Psi$. Due to the definition of $\Psi_{R,u}$, the valid 1-types of $\Psi_{R,u}$ must have the form $C_{i}(x)\wedge C^{A}_{i^{\prime}}(x)$, where $C^{A}_{i^{\prime}}(x)$ is defined as
@@
$r_{(i,i^{\prime}),(j,j^{\prime})}=\begin{cases}r_{i,j}=\mathsf{WMC}(\psi_{i,j}(a,b),w,\overline{w}),&i^{\prime}=j^{\prime},\\ \begin{aligned} r^{\not-}_{i,j}=\mathsf{WMC}(&\psi_{i,j}(a,b)\wedge\lnot R(a,b)\wedge\lnot R(b,a),w,\overline{w}),\end{aligned}&i^{\prime}\neq j^{\prime},\\ \end{cases}$
@@
Figure 1: The 1-type structure of $\Psi_{R,u}$. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replica of the original 1-types in each layer. The DP algorithm proceeds from left (cells $0,\cdots,3$) to right (cell $4$).
@@
$c_{(1,1)},\cdots,c_{(L,1)},\ c_{(1,2)},\cdots,c_{(L,2)},\ \cdots,\ c_{(1,u)},\cdots,c_{(L,u)}$
where $c_{(i,i^{\prime})}$ is the number of elements in the 1-type $(i,i^{\prime})$. For any $\hat{u}\in[u]$, we define $h_{\hat{u},(c_{1},c_{2},\dots,c_{L})}$ to be the weights of 1-types $(1,1),\cdots,(L,1),\cdots,(1,\hat{u}),\cdots,(L,\hat{u})$ such that $c_{i}=\sum_{j=1}^{\hat{u}}c_{(i,j)}$ for each $i\in[L]$, that is,
$h_{\hat{u},(c_{1},c_{2},\dots,c_{L})}$ $=\sum_{c_{(1,1)}+\cdots+c_{(1,\hat{u})}=c_{1}}\dots\sum_{c_{(L,1)}+\cdots+c_{(L,\hat{u})}=c_{L}}\frac{(\sum_{i\in[L]}c_{i})!}{\prod_{i\in[L],i^{\prime}\in[\hat{u}]}c_{(i,i^{\prime})}!}$
$\prod_{i\in[L],i^{\prime}\in[\hat{u}]}W(C_{i},w,\overline{w})^{c_{(i,i^{\prime})}}\cdot s_{i}^{\binom{c_{(i,i^{\prime})}}{2}}\prod_{\begin{subarray}{c}i,j\in[L],i^{\prime},j^{\prime}\in[\hat{u}],\\ (i,i^{\prime})<(j,j^{\prime})\end{subarray}}\left(r_{(i,i^{\prime}),(j,j^{\prime})}\right)^{c_{(i,i^{\prime})}c_{(j,j^{\prime})}},$
@@
For simplicity, we use $\bm{c}$ to denote $(c_{1},c_{2},\dots,c_{L})$. Then the new term $h_{\hat{u},\bm{c}}$ can be also computed by a DP algorithm:
@@
where recall that $\bm{\bar{c}}+\bm{c}^{*}=\bm{c}$ means that $\bar{c}_{i}+c^{*}_{i}=c_{i}$ for each $i\in[L]$,
$w_{in}(\bm{c}^{*})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{i=1}^{L}W(C_{i},w,\overline{w})^{c_{i}^{*}}\cdot s_{i}^{\binom{c_{i}^{*}}{2}}\prod_{1\leq i<j\leq L}\left(r_{i,j}\right)^{c_{i}^{*}c_{j}^{*}}$
@@
$w_{cross}(\bm{\bar{c}},\bm{c}^{*})=\prod_{i=1}^{L}\prod_{j=1}^{L}\left(r_{i,j}^{\not-}\right)^{\bar{c}_{i}c_{j}^{*}}$
@@
The new DP algorithm is summarized in Algorithm 1, where $\dots$) takes $n+1$ points as input and returns the corresponding polynomial of degree $n$.
@@ 3.3 Applications
In this section, we show how WCP can be used for computing WFOMC of $\text{C}^{2}$ sentences with cardinality constraints and several combinatorial axioms in polynomial time, i.e. we show that the fragments together with each of these axioms are domain-liftable. Some of these axioms were previously shown to be domain-liftable (e.g., connectedness axioms [MBS23], tree axiom [vBK23] and forest axiom [MBS23]), while for others we show their domain-liftability for the first time (e.g., generalized connectedness axiom and bipartite axiom). We also show that WCP allows us to efficiently compute the Tutte polynomial of any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence, e.g., block-structured graphs.
@@ 3.3.1 Axioms: Existing, Generalized, and New
$k$-connected-component axiom. As in Proposition 3.2, WCP associates the weight of each model with a term indicating the number of connected components in $G(R)$, therefore we can restrict the number of connected components of $G(R)$. The $k$-connected-component axiom to the binary predicate $R$, denoted by $connected_{k}(R)$, requires $R$ to be symmetric and $G(R)$ to consist of exactly $k$ connected components. As a special case, $connected_{1}(R)$, which represents that $G(R)$ is an undirected connected graph, i.e., has just one connected component, was already shown to be domain-liftable in [MBS23]. Here we generalize their result to $connected_{k}(R)$.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, the weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. By Proposition 3.2, we have that $\mathsf{WFOMC}(\Psi\land connected_{k}(R),n,w,\overline{w})$ is equal to $[u^{k}]f_{n}(u-1;\Psi_{R},w,\overline{w},R)$. ∎
Bipartite axiom. The bipartite axiom to the binary predicate $R$, denoted by $bipartite(R)$, requires $R$ to be symmetric and $G(R)$ to be a bipartite graph888 Note that bipartite graphs are not finitely axiomatizable in first-order logic by the compactness theorem..
@@
Tree axiom. The tree axiom [vBK23] to the binary predicate $R$ (denoted by $tree(R)$) requires $R$ to be symmetric and $G(R)$ to be a tree. To handle the tree axiom, we extend the ability of WCP to classify the models of a sentence by both the number of connected components in $G(R)$ and the number of edges in $G(R)$ by introducing a new variable to the weight of $R$.
@@
$f_{n}(u,v;\Psi,w,\overline{w},R)=\mathsf{WFOMC}(\Psi_{R,u},n,w_{R,v},\overline{w}),$
@@
Corollary 3.9 (Corollary from Proposition 3.2).
@@
$f_{n}(u,v;\Psi,w,\overline{w},R)=\sum_{\mu\in\mathcal{M}_{\Psi,n}}W(\mu,w,\overline{w})\cdot(u+1)^{cc(\mu_{R})}v^{2e(\mu_{R})},$
where $cc(\mu_{R})$ is the number of connected components of $G(\mu_{R})$ and $e(\mu_{R})$ is the number of undirected edges [^9] Here, as $R$ is symmetric, $G(\mu_{R})$ is interpreted as an undirected graph, and the number of undirected edges is equal to half the cardinality of $R$ in $\mu$. It is where the number $2$ in the exponent comes from. of $G(\mu_{R})$.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the tree axiom requires that $G(R)$ should be connected and the number of edges in $G(R)$ should be $n-1$. By Corollary 3.9, we have
$\mathsf{WFOMC}(\Psi\land$ $tree(R),n,w,\overline{w})=[uv^{2(n-1)}]f_{n}(u-1,v;\Psi_{R},w,\overline{w},R).$
Forest axiom. The forest axiom [MBS23] to the binary predicate $R$ (denoted by $forest(R)$) requires $R$ to be symmetric and $G(R)$ to be a forest. Similar to the tree axiom, we can also represent that the forest axiom is domain-liftable by Corollary 3.9.
@@
$\mathsf{WFOMC}($ $\Psi\land forest(R),n,w,\overline{w})=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]f_{n}(u-1,v;\Psi_{R},w,\overline{w},R).$
@@
The tree axiom and the forest axiom have already been shown to be domain-liftable in [vBK23] and [MBS23] respectively. Nevertheless, our method provides a more flexible command of the graph structures. For example, it is easy to add further constraints on the number of trees in the forest by our method (we will discuss this in Section 5).
@@
In a social network, one may want to model the compactness in terms of a symmetric relation $friends$ by the number of connected components in $G(friends)$. In this case, a real number $d$ might be used to represent the degree of compactness: the larger $d$ means that the network is more compact, i.e., the number of connected components in $G(friends)$ is smaller. This can be achieved by defining the weight of a possible world $\omega$ as $W(\omega,w,\overline{w})\cdot\exp(-d\cdot cc(\omega_{friends}))$. Then the reduced WFOMC of the inference problem is equal to $f_{n}(\exp(-d)-1;\Psi,w,\overline{w},friends)$, where $\Psi$ and $(w,\overline{w})$ are the sentence and weighting functions modeling other aspects of the social network.

### v3

@@ 2.5 Axioms
While $\text{C}^{2}$ captures numerous graph structures, it falls short in expressing certain important graph properties, such as being a tree, a forest, and strongly connected. In fact, these structures cannot be finitely axiomatized by first-order logic [libkin2004elements].
@@
An axiom is a special constraint on the interpretation of a binary predicate $R$ that $G(R)$ should be a certain combinatorial structure. We often write an axiom as $axiom(R)$ where axiomaxiom is an identifier of the axiom. For example, in what follows, we will discuss the tree axiom $tree(R)$, which requires $G(R)$ to be a tree, and the forest axiom $forest(R)$, which requires $G(R)$ to be a forest. For better illustration, we treat axioms as atomic formulas in a sentence, and use $\Psi\land axiom(R)$ to denote the sentence $\Psi$ with the axiom $axiom(R)$ conjoined. For instance, $tree(E)\land\forall x\left(Leaf(x)\leftrightarrow\exists_{=1}y\ E(x,y)\right)$ is a sentence that expresses the concept of trees with a leaf predicate $Leaf/1$.
@@ 2.6 Graph Polynomial
Tutte Polynomial. The Tutte polynomial [tuttepoly] is one of the most well-known graph polynomials for undirected graphs. Consider an undirected graph $G=(V,E)$, where $V$ is the set of vertices and $E$ is the set of edges. The Tutte polynomial for the graph $G$ is a polynomial in two variables $x$ and $y$ defined as
@@
The Tutte polynomial holds significant combinatorial meaning. Evaluating the polynomial at specific points reveals various combinatorial properties. For example, $T_{G}(1,1)$ counts the number of spanning forests of $G$ (specifically, the number of spanning trees if $G$ is connected), and $T_{G}(0,2)$ counts the strongly connected orientations of $G$ [tuttepoly-20]. Substituting a concrete value for only one variable yields another graph polynomial of $G$. For example, $T_{G}(x,0)$ corresponds to the chromatic polynomial [read1968introduction] of $G$; $T_{G}(0,y)$ corresponds to the flow polynomial [diestel2008graph] of $G$.
The computation of the Tutte polynomial for an arbitrary graph is known to be computationally challenging. Even the evaluation of a specific point of the polynomial is known to be $\mathsf{\#P}$-hard, with only a limited number of points being computationally tractable [tuttepoly-hardness].
Directed Chromatic Polynomial. The idea of the chromatic polynomial of a directed graph appeared in [chromatic-directed], although it was originally stated in terms of ordered sets. Given a directed graph $D=(V,E)$ where $n=|V|$, let $\chi_{D}(x)$ be the number of ways of coloring each vertex in $D$ by one of the colors $\{1,2,\cdots,x\}$ such that if there is an edge from $u$ to $v$, the color of $u$ is smaller than the color of $v$. We have the following lemma to state that $\chi_{D}(x)$ is a polynomial of $x$.
Lemma 2.13 ([chromatic-directed]).
@@
Lemma 2.16 ([chromatic-directed, tutte-directed]).
@@
[…]
@@ 3.1 Definition
Let $\Psi$ be a first-order logic sentence possibly with cardinality constraints and ground unary literals, $w,\overline{w}$ be two weighting functions, and $R$ be a distinguished binary relation. The $n$-th WCP $f_{n}(u)$ of $\Psi$ for the relation $R$ is the univariate polynomial which satisfies:
@@
For the sentence $\Psi$, the weighting functions $w$ and $\overline{w}$ and the binary relation $R$, it holds that
@@
The $n$-th WCP $f_{n}(u;\Psi,w,\overline{w},R)$ always exists, and it is a unique univariate polynomial of degree at most $n$.
@@
The existence and the upper bound of the degree follow immediately from Proposition 3.2 since $cc(\mu_{R})\leq n$. ∎
@@ 3.2 Fast Calculation
Note that by definition, evaluating the polynomial at a specific positive integer is a WFOMC query. However, such WFOMC is not readily solvable by existing algorithms (e.g., [WFOMC-FO3, WFOMC-FO2-faster]) even for the $\text{FO}^{2}$ fragment. In general, the runtime of these algorithms is exponential in the length of the sentence, while the length of the sentence $\Psi_{R,u}$ in the WFOMC query can be as large as the domain size $n$, hence these algorithms would not lead to polynomial runtime in the domain size. In particular, the algorithm described in Section 2.4 is not sufficient in this case either, as there would be $O(2^{n})$ 1-types and, consequently, an exponential time data complexity. To this end, we adapt this algorithm into a dynamic programming (DP) style algorithm that exploits the structure of $\Psi_{R,u}$ to achieve an efficient calculation of the polynomials for $\text{FO}^{2}$ sentences, which then applies to $\text{C}^{2}$ sentences with cardinality constraints by techniques mentioned in Section 2.4.
Let $\Psi$ and $R$ be the input $\text{FO}^{2}$ sentence and the binary relation that WCP is defined on. Consider the point evaluation of the WCP at $u$. Let $C_{1}(x),C_{2}(x),\cdots,\linebreak C_{p}(x)$ be the valid 1-types  [^7] We call a 1-type valid if it is satisfiable in some model of the sentence. of $\Psi$. Due to the definition of $\Psi_{R,u}$, the valid 1-types of $\Psi_{R,u}$ must have the form $C_{i}(x)\wedge C^{A}_{i^{\prime}}(x)$, where $C^{A}_{i^{\prime}}(x)$ is defined as
@@
[…]
@@
Figure 1: The 1-type structure of $\Psi_{R,u}$. Each tiny gray box represents an original 1-type of $\Psi$. The bold gray cells (each containing $4\times 4$ boxes) refer to the replica of the original 1-types in each layer. The DP algorithm proceeds from left (cells $0,\cdots,3$) to right (cell 44).
@@
$c_{(1,1)},\cdots,c_{(p,1)},\ c_{(1,2)},\cdots,c_{(p,2)},\ \cdots,\ c_{(1,u)},\cdots,c_{(p,u)}$
[…]
$h_{\hat{u},(c_{1},c_{2},\dots,c_{p})}$ $=\sum_{c_{(1,1)}+\cdots+c_{(1,\hat{u})}=c_{1}}\dots\sum_{c_{(p,1)}+\cdots+c_{(p,\hat{u})}=c_{p}}\frac{(\sum_{i\in[p]}c_{i})!}{\prod_{i\in[p],i^{\prime}\in[\hat{u}]}c_{(i,i^{\prime})}!}$
$\prod_{i\in[p],i^{\prime}\in[\hat{u}]}W(C_{i},w,\overline{w})^{c_{(i,i^{\prime})}}\cdot s_{i}^{\binom{c_{(i,i^{\prime})}}{2}}\prod_{\begin{subarray}{c}i,j\in[p],i^{\prime},j^{\prime}\in[\hat{u}],\\ (i,i^{\prime})<(j,j^{\prime})\end{subarray}}\left(r_{(i,i^{\prime}),(j,j^{\prime})}\right)^{c_{(i,i^{\prime})}c_{(j,j^{\prime})}},$
@@
For simplicity, we use $\bm{c}$ to denote $(c_{1},c_{2},\dots,c_{p})$. Then the new term $h_{\hat{u},\bm{c}}$ can be also computed by a DP algorithm:
@@
where recall that $\bm{\bar{c}}+\bm{c}^{*}=\bm{c}$ means that $\bar{c}_{i}+c^{*}_{i}=c_{i}$ for each $i\in[p]$,
$w_{in}(\bm{c}^{*})=\binom{|\bm{c}^{*}|}{\bm{c}^{*}}\prod_{i=1}^{p}W(C_{i},w,\overline{w})^{c_{i}^{*}}\cdot s_{i}^{\binom{c_{i}^{*}}{2}}\prod_{1\leq i<j\leq p}\left(r_{i,j}\right)^{c_{i}^{*}c_{j}^{*}}$
@@
$w_{cross}(\bm{\bar{c}},\bm{c}^{*})=\prod_{i=1}^{p}\prod_{j=1}^{p}\left(r_{i,j}^{\not-}\right)^{\bar{c}_{i}c_{j}^{*}}$
@@
The new DP algorithm is summarized in Algorithm 1, where interpolate_1d($\dots$) takes $n+1$ points as input and returns the corresponding polynomial of degree $n$.
@@ 3.3 Applications
In this section, we show how WCP can be used for computing WFOMC of $\text{C}^{2}$ sentences with cardinality constraints and several combinatorial axioms in polynomial time, i.e. we show that the fragments together with each of these axioms are domain-liftable. Some of these axioms were previously shown to be domain-liftable (e.g., connectedness axioms [WFOMC-axioms], tree axiom [WFOMC-tree-axioms] and forest axiom [WFOMC-axioms]), while for others we show their domain-liftability for the first time (e.g., generalized connectedness axiom and bipartite axiom). We also show that WCP allows us to efficiently compute the Tutte polynomial of any graph that can be encoded by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence, e.g., block-structured graphs.
@@ 3.3.1 Axioms: Existing, Generalized, and New
$k$-connected-component axiom. As in Proposition 3.2, WCP associates the weight of each model with a term indicating the number of connected components in $G(R)$, therefore we can restrict the number of connected components of $G(R)$. The $k$-connected-component axiom to the binary predicate $R$, denoted by $connected_{k}(R)$, requires $R$ to be symmetric and $G(R)$ to consist of exactly $k$ connected components. As a special case, $connected_{1}(R)$, which represents that $G(R)$ is an undirected connected graph, i.e., has just one connected component, was already shown to be domain-liftable in [WFOMC-axioms]. Here we generalize their result to $connected_{k}(R)$.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, the weighting functions $w,\overline{w}$, and a distinguished binary predicate $R$. By Proposition 3.2, we have that $\mathsf{WFOMC}(\Psi\land connected_{k}(R),n,w,\overline{w})$ is equal to $u^{k}]f_{n}(u-1;\Psi_{R},w,\overline{w},R)$. ∎
Bipartite axiom. The bipartite axiom to the binary predicate $R$, denoted by $bipartite(R)$, requires $R$ to be symmetric and $G(R)$ to be a bipartite graph [^8] Note that bipartite graphs are not finitely axiomatizable in first-order logic by the compactness theorem..
@@
Tree and forest axioms. The tree axiom [WFOMC-tree-axioms] to the binary predicate $R$ (denoted by $tree(R)$) requires $R$ to be symmetric and $G(R)$ to be a tree. The forest axiom [WFOMC-axioms] to the binary predicate $R$ (denoted by $forest(R)$) requires $R$ to be symmetric and $G(R)$ to be a forest. To handle these two axioms, we extend the ability of WCP to classify the models of a sentence by both the number of connected components in $G(R)$ and the number of edges in $G(R)$ by introducing a new variable to the weight of $R$.
@@
$\hat{f}_{n}(u,v;\Psi,w,\overline{w},R)=\mathsf{WFOMC}(\Psi_{R,u},n,w_{R,v},\overline{w}),$
@@
Corollary 3.9 (Corollary from Propositions 3.2 and 3.5).
@@
$\hat{f}_{n}(u,v;\Psi,w,\overline{w},R)=\sum_{\mu\in\mathcal{M}_{\Psi,n}}W(\mu,w,\overline{w})\cdot(u+1)^{cc(\mu_{R})}v^{2e(\mu_{R})},$
where $cc(\mu_{R})$ is the number of connected components of $G(\mu_{R})$ and $e(\mu_{R})$ is the number of undirected edges [^9] Here, as $R$ is symmetric, $G(\mu_{R})$ is interpreted as an undirected graph, and the number of undirected edges is equal to half the cardinality of $R$ in $\mu$. It is where the number $2$ in the exponent comes from.of $G(\mu_{R})$. Furthermore, the extended WCP of $\Psi$ always exists, and can be computed in time polynomial in $n$ using the same techniques as in Proposition 3.5.
@@
Consider a $\text{C}^{2}$ sentence $\Psi$ with cardinality constraints, a domain of size $n$, weighting functions $w,\overline{w}$ and a distinguished binary predicate $R$. By definition, the tree axiom requires that $G(R)$ should be connected and the number of edges in $G(R)$ should be n1n-1. By Corollary 3.9, we have
$\mathsf{WFOMC}(\Psi\land$ $tree(R),n,w,\overline{w})=[uv^{2(n-1)}]\hat{f}_{n}(u-1,v;\Psi_{R},w,\overline{w},R).$
@@
$\mathsf{WFOMC}($ $\Psi\land forest(R),n,w,\overline{w})=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]\hat{f}_{n}(u-1,v;\Psi_{R},w,\overline{w},R).$
@@
The tree axiom and the forest axiom have already been shown to be domain-liftable in [WFOMC-tree-axioms] and [WFOMC-axioms] respectively. Nevertheless, our method provides a more flexible command of the graph structures. For example, it is easy to add further constraints on the number of trees in the forest by our method (we will discuss this in Section 5).
@@
In a social network, one may want to model the compactness in terms of a symmetric relation friendsfriends by the number of connected components in $G(friends)$. In this case, a real number $d$ might be used to represent the degree of compactness: the larger $d$ means that the network is more compact, i.e., the number of connected components in $G(friends)$ is smaller. This can be achieved by defining the weight of a possible world $\omega$ as $W(\omega,w,\overline{w})\cdot\exp(-d\cdot cc(\omega_{friends}))$. Then the reduced WFOMC of the inference problem is equal to $f_{n}(\exp(-d)-1;\Psi,w,\overline{w},friends)$, where $\Psi$ and $(w,\overline{w})$ are the sentence and weighting functions modeling other aspects of the social network.